xx IllinoisGRMHD

IllinoisGRMHD

Samuel Cupp Leonardo Rosa Werneck Terrence Pierre Jacques

Zachariah Etienne

Date

Abstract

IllinoisGRMHD solves the equations of general relativistic magneto-hydrodynamics (GRMHD) using the General Relativistic Hydrodynamics Library (GRHayL), which is included in the Einstein Toolkit via the GRHayLib thorn. The core library inherits or adapts code from the original IllinoisGRMHD thorn into a modular, infrastructure-agnostic library. This original thorn is itself a rewrite of the Illinois Numerical Relativity (ILNR) group’s GRMHD code (ca. 2009–2014), which was used in their modeling of the following systems:

Magnetized circumbinary disk accretion onto binary black holes

Magnetized black hole–neutron star mergers

Magnetized Bondi flow, Bondi-Hoyle-Littleton accretion White dwarf–neutron star mergers

The improved GRHayL-based IllinoisGRMHD incorporates more options for controlling the evolution, most notably the addition of new conservative-to-primitive solvers and nuclear equations of state.

1 Introduction

This thorn provides a GRMHD evolution code built on the General Relativistic Hydrodynamics Library (GRHayL), which is included in the Einstein Toolkit via the GRHayLib thorn. The core library inherits or adapts code from the original IllinoisGRMHD thorn into a modular, infrastructure-agnostic library.

The new and improved IllinoisGRMHD provides support for hybrid and tabulated EOS, both with and without entropy. The core functions all come from GRHayLib, ensuring that any improvements or updates to the library can be easily adopted by the thorn. The conservative-to-primitive solver is selected at runtime, allowing for users to take advantage of any Con2Prim methods provided by

GRHayL.

2 Parameters

Most of the behavior of this thorn is controlled by GRHayLib, the thorn that provides GRHayL functionality within the Einstein Toolkit. However, there are several parameters in the thorn which control diagnostic or debugging features. Most are self-explanatory, but some parameters deserve a more detailed explanation.

First, many analysis or diagnostic thorns use HydroBase variables. By default, IllinoisGRMHD never copies the data back from IllinoisGRMHD variables to HydroBase variables, which would prevent the usage of these thorns. The ConvertxtoxHydroBasexevery parameter sets how often to copy back this data. Naturally, copying more frequently will slow down the simulation, so this should be set on the same frequency as the analysis or IO thorn that is using the HydroBase variable. We reduced the amount of data copying in IllinoisGRMHD by using the HydroBase variables directly, which removes the need to copy them back and removes extra grid functions from memory. We still use distinct velocity variables because we use a different velocity than is defined in

HydroBase.

Next, there are several parameters for debugging or other testing. The perturbing parameters give users the ability to affect either the initial data or evolution data at runtime without slowing down standard simulations. The perturbxinitialxdata parameter controls whether the initial data from HydroBase is perturbed, and the perturbxeveryxcon2prim parameter controls whether the conservative variables are perturbed before every con2prim call. The magnitude of these perturbations is set by randomxpert.

3 Updating Old Parfiles

With the many changes to improve the code and transition to using GRHayLib, parfile setups have changed significantly. To this end, we provide a guide to updating parfiles from old IllinoisGRMHD to the GRHayL-based IllinoisGRMHD. Old parfiles are temporarily supported via backward compatibility patch, but this support ends in the ETx2024x11 release.

First, the two thorns IDxconverterxILGRMHD and ConvertxtoxHydroBase are entirely deprecated and will be removed. As such, these thorns and all their parameters should be removed from the parfile. Their equivalent parameters are as follows:

11  Old  11  New
11 IDxconverterxILGRMHD::GammaxInitial  11 GRHayLib::Gammaxppolyxin[0]
11 IDxconverterxILGRMHD::randomxseed  11 IllinoisGRMHD::randomxseed
11 IDxconverterxILGRMHD::randomxpert  11 IllinoisGRMHD::randomxpert
11 IDxconverterxILGRMHD::KxInitial  11 GRHayLib::kxppoly0
11 ConvertxtoxHydroBase::ConvertxtoxHydroBasexevery 11 IllinoisGRMHD::ConvertxtoxHydroBasexevery

11

The parameter IDxconverterxILGRMHD::purexhydroxrun is entirely deprecated, as the GRHayLHD thorn provides this functionaility. Additionally, to properly trigger the random perturbation the new parameter IllinoisGRMHD::perturbxinitialxdata should be set.

The parameters in IllinoisGRMHD have also changed significantly, with many moving to the GRHayLib thorn. These changes are as follows:

11  Old  11  New
11 IllinoisGRMHD::GAMMAxSPEEDxLIMIT 11 GRHayLib::maxxLorentzxfactor
11 IllinoisGRMHD::Kxpoly  11 GRHayLib::kxppoly0
11 IllinoisGRMHD::rhoxbxatm  11 GRHayLib::rhoxbxatm
11 IllinoisGRMHD::rhoxbxmax  11 GRHayLib::rhoxbxmax
11 IllinoisGRMHD::Psi6threshold  11 GRHayLib::Psi6threshold
11 IllinoisGRMHD::neos  11 GRHayLib::neos
11 IllinoisGRMHD::gammaxth  11 GRHayLib::Gammaxth
11 IllinoisGRMHD::dampxlorenz  11 GRHayLib::Lorenzxdampingxfactor

11

As seen here, the ID converter and IllinoisGRMHD thorns both had a parameter. These are condensed into a single parameter in GRHayLib.

Two parameters are deprecated and marked for removal. These are IllinoisGRMHD::tauxatm and IllinoisGRMHD::conservxtoxprimsxdebug. The former is now automatically computed by GRHayLib, and the latter’s feature is no longer available.

Other IllinoisGRMHD parameter names are unchanged. However, some of their options are deprecated and also marked for removal. These are the ”essential” and ”essential+iteration output” options for IllinoisGRMHD::verbose.

Finally, note that the magnetic field quantities have been rescaled by , as this simplifies many equations in GRHayL. This definition is consistent with the magnetic field defined by HydroBase and GRHydro (which the old IllinoisGRMHD did not respect). To maintain support for old initial data thorns, we by default assume that the HydroBase , use the old definition and rescale them when copying into the IllinoisGRMHD variables. The parameter rescalexmagnetics controls this behavior. If the thorns using the old definition (i.e. thorns in wvuthorns and wvuthornsxdiagnostics) are updated to use the new definition, this parameter will be switch such that the default behaves as expected for these thorns. Such a change will be detailed in the Einstein Toolkit release notes prior to the release introducing the change.

In addition to these changes, GRHayLib offers new parameter options for controlling the simulation, such as a wider range of conservative-to-primitive routines, new equation of state options, and control over the PPM reconstruction parameters. Please see that thorn’s documentation for more information about these features, as well as GRHayL’s documentation on the library’s implementation details.

4 Acknowledgements

Note that many codes have fed into the GRHayL library. The most prominent is the IllinoisGRMHD thorn developed primarily by Zachariah Etienne, Yuk Tung Liu, and Vasileios Paschalidisx[1]. This thorn is based on the GRMHD code of the Illinois Numerical Relativity group (ca. 2014), written by Matt Duez, Yuk Tung Liu, and Branson Stephensx.

For the full list of contributions and related citations, please use GRHayL’s citation guide.

References

[1]xxxZ.xB.xEtienne, V.xPaschalidis, R.xHaas, P.xMxsta and S.xL.xShapiro, “IllinoisGRMHD: An Open-Source, User-Friendly GRMHD Code for Dynamical Spacetimes,” Class. Quant. Grav. 32, 175009 (2015) doi:10.1088/0264-9381/32/17/175009 [arXiv:1501.07276 [astro-ph.HE]].

5 Parameters

11 conservxtoxprimsxdebug 10 Scope: private01  INT

11 Description: Deprecated.

11 Range 10 01 Default: (none)
11 0:1  10 zero (no) or one (yes)

11

11 convertxtoxhydrobasexevery 10 Scope: private01  INT

11 Description: How often to convert IllinoisGRMHD primitive variables to HydroBase (Valencia formulation) primitive variables? Needed for some ET-based diagnostics. NOT needed for pure IllinoisGRMHD runs.

11  Range  10              01  Default: (none)
11 0:*  10 zero (disable) or positive (every N iterations)

11

11 dampxlorenz  10 Scope: private01  REAL

11 Description: Deprecated. Please use GRHayLib.

11 Range 10 01 Default: 0.0
11 *:*  10 any real

11

11 emxbc  10 Scope: private01  KEYWORD

11 Description: EM field boundary condition

11 Range 10 01 Default: copy
11 copy  10 Copy data from nearest boundary point
11 frozen  10 Frozen boundaries

11

11 gammaxspeedxlimit  10 Scope: private01  REAL

11 Description: Deprecated. Please use GRHayLib.

11 Range 10 01 Default: 10.0
11 1:*  10 Positive > 1, though you’ll likely have troubles far above 10.

11 gammaxth  10 Scope: private01  REAL

11 Description: Deprecated. Please use GRHayLib.

11 Range 10 01 Default: -1
11 0:*  10 Physical values
11 -1  10 forbidden value to make sure it is explicitly set in the parfile

11

11 kxpoly  10 Scope: private01  REAL

11 Description: Deprecated. Please use GRHayLib.

11 Range 10 01 Default: 1.0
11 0:*  10 Positive

11

11 matterxbc  10 Scope: private01  KEYWORD

11 Description: Chosen Matter boundary condition

11 Range 10 01 Default: outflow
11 copy  10 Copy BCs
11 outflow  10 Copy BCs with additional outflow BC for velocities
11 frozen  10 Frozen boundaries

11

11 neos  10 Scope: private01  INT

11 Description: Deprecated. Please use GRHayLib.

11 Range 10 01 Default: 1
11 1:10  10 Any integer between 1 and 10

11

11 perturbxeveryxcon2prim 10 Scope: private01  BOOLEAN

11 Description: Random perturbation applied every time con2prim runs

11                       10              01  Default: no

11

11 perturbxinitialxdata  10 Scope: private01  BOOLEAN

11 Description: Random perturbation to initial data

11                       10              01  Default: no

11

11 psi6threshold  10 Scope: private01  REAL

11 Description: Deprecated. Please use GRHayLib.

11 Range 10 01 Default: 1e100
11 *:*  10 Can set to anything

11

11 randomxpert  10 Scope: private01  REAL

11 Description: Magnitude of random perturbations

11 Range 10 01 Default: (none)
11 *:*  10 Anything goes.

11 randomxseed  10 Scope: private01  INT

11 Description: Random seed for random, generally roundoff-level perturbation on initial data. Seeds srand(), and rand() is used for the RNG.

11  Range  10              01  Default: (none)
11 0:99999999  10 Anything unsigned goes.

11

11 rescalexmagnetics  10 Scope: private01  BOOLEAN

11 Description: Rescale magnetic data to match original IllinoisGRMHD definitions

11                       10              01  Default: yes

11

11 rhoxbxatm  10 Scope: private01  REAL

11 Description: Deprecated. Please use GRHayLib.

11 Range 10 01 Default: 1e200
11 *:*  10 Allow for negative values. This enables us to debug the code and verify if rhoxbxatm is properly set.

11

11 rhoxbxmax  10 Scope: private01  REAL

11 Description: Deprecated. Please use GRHayLib.

11 Range 10 01 Default: 1e300
11 0:*  10 Note that you will have problems unless rhoxbxatm<rhoxbxmax

11

11 symxbz  10 Scope: private01  REAL

11 Description: In-progress equatorial symmetry support: Symmetry parameter across z axis for magnetic fields = +/- 1

11  Range  10              01  Default: 1.0
11 -1.0:1.0  10 Set to +1 or -1.

11

11 symmetry  10 Scope: private01  KEYWORD

11 Description: Currently only no symmetry supported, though work has begun in adding equatorial-symmetry support.

11  Range  10              01  Default: none
11 none  10 no symmetry, full 3d domain

11

11 tauxatm  10 Scope: private01  REAL

11 Description: Deprecated. Please use GRHayLib.

11 Range 10 01 Default: 1e100
11 0:*  10 Positive

11

11 updatextmunu  10 Scope: private01  BOOLEAN

11 Description: Update Tmunu, for RHS of Einstein’s equations?

11                       10              01  Default: yes

11 verbose  10 Scope: private01  KEYWORD

11 Description: Determines how much evolution information is output

11 Range 10 01 Default: yes
11 no  10 Complete silence
11 yes  10 ”Outputs information about conservative-to-prim itive fixes, failures, etc.”
11 essential  10 Deprecated option.
11 see [1] below  10 Deprecated option.

11

essential+iterationxoutput

11 eosxtype  10 Scope: shared from GRHAYLIB01KEYWORD

11

11 evolvexentropy  10 Scope: shared from GRHAYLIB01BOOLEAN

11

6 Interfaces

General Implements: illinoisgrmhd

Inherits: admbase tmunubase hydrobase grhaylib

Grid Variables

6.0.1 PRIVATE GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 grmhdxvelocities 10                    01  compact 11 0
11                    10  vx  01 description 11 Components of primitive three velocity vx. Note that vx is defined in terms of 4-velocity as: vx = ux/uˆ0 . Note that this definition differs from the Valencia formalism.
11                    10  vy  01 dimensions 11 3
11                    10  vz  01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 InterpNumTimelevels=1 prolongation=”none”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 u0 10 u0 01 compact 11 0
11                    10                    01 description 11 4-velocity component uˆ0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 InterpNumTimelevels=1 prolongation=”none” checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 grmhdxconservatives 10                    01  compact 11 0
11                    10  rhoxstar  01 description 11 Evolved hydrodynamic variables
11                    10  tau  01 dimensions 11 3
11                    10  Stildex  01 distribution 11 DEFAULT
11                    10  Stildey  01 group type 11 GF
11                    10  Stildez  01  tags 11 prolongation=”ENO”
11                    10                    01 timelevels 11 3
11                    10                    01 variable type 11 REAL
11 entxstar 10 entxstar 01 compact 11 0
11                    10                    01 description 11 Evolved entropy variable
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 prolongation=”ENO”
11                    10                    01 timelevels 11 3
11                    10                    01 variable type 11 REAL
11 yexstar 10 Yexstar 01 compact 11 0
11                    10                    01 description 11 Evolved electron fraction variable
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 prolongation=”ENO”
11                    10                    01 timelevels 11 3
11                    10                    01 variable type 11 REAL
11 ax 10 Ax 01 compact 11 0
11                    10                    01 description 11 x-component of the vector potential
11 10 x 01 description 11 evolved when constrainedxtransportxscheme==3
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 Prolongation=”STAGGER011”
11                    10                    01 timelevels 11 3
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 ay 10 Ay 01 compact 11 0
11                       10                    01 description 11 y-component of the vector potential
11 10 x 01 description 11 evolved when constrainedxtransportxscheme==3
11                       10                    01 dimensions 11 3
11                       10                    01 distribution 11 DEFAULT
11                       10                    01 group type 11 GF
11                       10                    01  tags 11 Prolongation=”STAGGER101”
11                       10                    01 timelevels 11 3
11                       10                    01 variable type 11 REAL
11 az 10 Az 01 compact 11 0
11                       10                    01 description 11 z-component of the vector potential
11 10 x 01 description 11 evolved when constrainedxtransportxscheme==3
11                       10                    01 dimensions 11 3
11                       10                    01 distribution 11 DEFAULT
11                       10                    01 group type 11 GF
11                       10                    01  tags 11 Prolongation=”STAGGER110”
11                       10                    01 timelevels 11 3
11                       10                    01 variable type 11 REAL
11 phitilde 10 phitilde 01 compact 11 0
11                       10                    01 description 11 sqrt{gamma} Phi
11 10 x 01 description 11 where Phi is the em scalar potential
11                       10                    01 dimensions 11 3
11                       10                    01 distribution 11 DEFAULT
11                       10                    01 group type 11 GF
11                       10                    01  tags 11 Prolongation=”STAGGER111”
11                       10                    01 timelevels 11 3
11                       10                    01 variable type 11 REAL
11  grmhdxbxstagger  10                    01  compact 11 0
11                       10  Bxxstagger  01 description 11 Densitized B-field components defined at staggered points [Bxxstagger at (i+1/2
11                       10  x  01 description 11 j
11                       10  Bxxstagger  01 description 11 k)
11                       10  Bxxstagger  01 description 11 Byxstagger at (i
11                       10  Bxxstagger  01 description 11 j+1/2
11                       10  Bxxstagger  01 description 11 k)
11                       10  Bxxstagger  01 description 11 Bzxstagger at (i
11                       10  Bxxstagger  01 description 11 j
11                       10  Bxxstagger  01 description 11 k+1/2)].
11                       10  Byxstagger  01 dimensions 11 3
11                       10  Bzxstagger  01 distribution 11 DEFAULT
11                       10                    01 group type 11 GF
11                       10                    01  tags 11 InterpNumTimelevels=1 prolongation=”none”
11                       10                    01 timelevels 11 1
11                       10                    01 variable type 11 REAL
11  grmhdxbxcenter  10                    01  compact 11 0
11                       10  Bxxcenter  01 description 11 B-field components defined at vertices.
11                       10  Byxcenter  01 dimensions 11 3
11                       10  Bzxcenter  01 distribution 11 DEFAULT
11                       10                    01 group type 11 GF
11                       10                    01  tags 11 InterpNumTimelevels=1 prolongation=”none”
11                       10                    01 timelevels 11 1
11                       10                    01 variable type 11 REAL
11 grmhdxconservativesxrhs 10                    01  compact 11 0
11                       10  rhoxstarxrhs  01 description 11 Right-hand side for evolved conservative variables
11                       10  tauxrhs  01 dimensions 11 3
11                       10  Stildexxrhs  01 distribution 11 DEFAULT
11                       10  Stildeyxrhs  01 group type 11 GF
11                       10  Stildezxrhs  01  tags 11 prolongation=”none” Checkpoint=”no”
11                       10                    01 timelevels 11 1
11                       10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 entxstarxrhs 10 entxstarxrhs 01 compact 11 0
11                        10                    01 description 11 Right-hand side for entxstar
11                        10                    01 dimensions 11 3
11                        10                    01 distribution 11 DEFAULT
11                        10                    01 group type 11 GF
11                        10                    01  tags 11 prolongation=”none” Checkpoint=”no”
11                        10                    01 timelevels 11 1
11                        10                    01 variable type 11 REAL
11 yexstarxrhs 10 Yexstarxrhs 01 compact 11 0
11                        10                    01 description 11 Right-hand side for Yexstar
11                        10                    01 dimensions 11 3
11                        10                    01 distribution 11 DEFAULT
11                        10                    01 group type 11 GF
11                        10                    01  tags 11 prolongation=”none” Checkpoint=”no”
11                        10                    01 timelevels 11 1
11                        10                    01 variable type 11 REAL
11  emxrhs  10                    01  compact 11 0
11                        10  Axxrhs  01 description 11 Storage for the right-hand side of the partialxt A equations and partialxt (psiˆ6 Phi) equation.
11                        10  Ayxrhs  01 dimensions 11 3
11                        10  Azxrhs  01 distribution 11 DEFAULT
11                        10  phitildexrhs  01 group type 11 GF
11                        10                    01  tags 11 prolongation=”none” Checkpoint=”no”
11                        10                    01 timelevels 11 1
11                        10                    01 variable type 11 REAL
11 grmhdxcminxcmaxxtemps 10                    01  compact 11 0
11                        10  cminxx  01 description 11 Store min and max characteristic speeds in all three directions.
11                        10  cmaxxx  01 dimensions 11 3
11                        10  cminxy  01 distribution 11 DEFAULT
11                        10  cmaxxy  01 group type 11 GF
11                        10  cminxz  01  tags 11 prolongation=”none” Checkpoint=”no”
11                        10  cmaxxz  01 timelevels 11 1
11                        10                    01 variable type 11 REAL
11 grmhdxfluxxtemps 10 01 compact 11 0
11                        10  rhoxstarxflux 01 description 11 Flux variables for computing RHSs
11                        10  tauxflux  01 dimensions 11 3
11                        10  Stildexxflux  01 distribution 11 DEFAULT
11                        10  Stildeyxflux  01 group type 11 GF
11                        10  Stildezxflux  01  tags 11 prolongation=”none” Checkpoint=”no”
11                        10                    01 timelevels 11 1
11                        10                    01 variable type 11 REAL
11 entxstarxflux 10 entxstarxflux 01 compact 11 0
11                        10                    01 description 11 Flux variable for entxstar
11                        10                    01 dimensions 11 3
11                        10                    01 distribution 11 DEFAULT
11                        10                    01 group type 11 GF
11                        10                    01  tags 11 prolongation=”none” Checkpoint=”no”
11                        10                    01 timelevels 11 1
11                        10                    01 variable type 11 REAL

11  x Group Names x  10 x Variable Names x01 Details x 11 x
11  yexstarxflux  10  Yexstarxflux  01  compact 11 0
11                                   10                    01 description 11 Flux variable for Yexstar
11                                   10                    01 dimensions 11 3
11                                   10                    01 distribution 11 DEFAULT
11                                   10                    01 group type 11 GF
11                                   10                    01  tags 11 prolongation=”none” Checkpoint=”no”
11                                   10                    01 timelevels 11 1
11                                   10                    01 variable type 11 REAL
11 grmhdxprimitivesxreconstructedxtemps 10                    01  compact 11 0
11                                   10  vxr  01 description 11 Temporary variables used for primitives reconstruction
11                                   10  vyr  01 dimensions 11 3
11                                   10  vzr  01 distribution 11 DEFAULT
11                                   10  vxl  01 group type 11 GF
11                                   10  vyl  01  tags 11 prolongation=”none” Checkpoint=”no”
11                                   10  vzl  01 timelevels 11 1
11                                   10  Bxxstaggerr  01 variable type 11 REAL
11  failurexchecker  10  failurexchecker 01  compact 11 0
11                                   10                    01 description 11 Gridfunction to track conservative-to-primitives solver fixes. Beware that this gridfunction is overwritten at each RK substep.
11                                   10                    01 dimensions 11 3
11                                   10                    01 distribution 11 DEFAULT
11                                   10                    01 group type 11 GF
11                                   10                    01  tags 11 prolongation=”none” Checkpoint=”no” InterpNumTimelevels=1
11                                   10                    01 timelevels 11 1
11                                   10                    01 variable type 11 REAL
11  emxpsi6phi  10                    01  compact 11 0
11                                   10  psi6phi  01 description 11 Deprecated variable
11                                   10                    01 dimensions 11 3
11                                   10                    01 distribution 11 DEFAULT
11                                   10                    01 group type 11 GF
11                                   10                    01  tags 11 Prolongation=”STAGGER111”
11                                   10                    01 timelevels 11 3
11                                   10                    01 variable type 11 REAL
11  grmhdxprimitivesxallbutbi  10                    01  compact 11 0
11                                   10  rhoxb  01 description 11 Deprecated variables
11                                   10  P  01 dimensions 11 3
11                                   10                    01 distribution 11 DEFAULT
11                                   10                    01 group type 11 GF
11                                   10                    01  tags 11 InterpNumTimelevels=1 prolongation=”none”
11                                   10                    01 timelevels 11 1
11                                   10                    01 variable type 11 REAL
11  grmhdxprimitivesxbi  10                    01  compact 11 0
11                                   10  Bx  01 description 11 Deprecated variables
11                                   10  By  01 dimensions 11 3
11                                   10  Bz  01 distribution 11 DEFAULT
11                                   10                    01 group type 11 GF
11                                   10                    01  tags 11 InterpNumTimelevels=1 prolongation=”none”
11                                   10                    01 timelevels 11 1
11                                   10                    01 variable type 11 REAL

11

Uses header: Symmetry.h GRHayLib.h

7 Schedule This section lists all the variables which are assigned storage by thorn GRHayLET/IllinoisGRMHD. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11 Always:  11 Conditional:
11 HydroBase::Bvec[1] HydroBase::Avec[1] HydroBase::Aphi[1]  11 Yexstar[3] Yexstarxrhs Yexstarxflux
11 grmhdxconservatives[3] Ax[3] Ay[3] Az[3] phitilde[3]  11 entxstar[3] entxstarxrhs entxstarxflux
11 grmhdxvelocities u0 grmhdxBxcenter grmhdxBxstagger  11 emxpsi6phi[3] grmhdxprimitivesxallbutBi grmhdxprimitivesxBi
11 grmhdxconservativesxrhs EMxrhs  11 x
11 grmhdxprimitivesxreconstructedxtemps grmhdxcminxcmaxxtemps failurexchecker 11 x
11 grmhdxfluxxtemps Yexstarxflux entxstarxflux  11 x

11 x  11 x

11

Scheduled Functions

MoLxRegister (conditional) x illinoisgrmhdxregistervars x register evolved, rhs variables in illinoisgrmhd for mol x

11 x 11 Language: 11 c
11 x 11 Options: 11 meta
11 x 11 Type:  11 function

11

DriverxBoundarySelect (conditional) x illinoisgrmhdxspecifyxdriverxbcs x register boundary conditions with presync x

11 x 11 Language: 11 c
11 x 11 Options: 11 meta
11 x 11 Type:  11 function

11

CCTKxPOSTPOSTINITIAL (conditional) x illinoisgrmhdxsetxgzxsymmetries x compute post-initialdata quantities

11 x 11 After:  11 con2prim
11 x 11 Language: 11 c
11 x 11 Reads:  11 grid::coordinates
11 x 11 x  11 grmhdxbxcenter
11 x 11 x  11 grmhdxbxstagger
11 x 11 x  11 phitilde

11 x 11 x  11 ax
11 x 11 x  11 ay
11 x 11 x  11 az

11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxbxcenter(boundary)
11 x 11 x  11 grmhdxbxstagger(boundary)
11 x 11 x  11 phitilde(boundary)
11 x 11 x  11 ax(boundary)
11 x 11 x  11 ay(boundary)
11 x 11 x  11 az(boundary)

11

HydroBasexCon2Prim (conditional) x illinoisgrmhdxcon2prim x compute primitive variables from conservatives x

11 x 11 Type: 11 group

11

IllinoisGRMHDxCon2Prim (conditional) x illinoisgrmhdxsync x sync conservative variable ghost zones for c2p routine x

11 x 11 Language: 11 c
11 x 11 Sync:  11 grmhdxconservatives
11 x 11 x  11 yexstar
11 x 11 x  11 entxstar

11 x 11 x  11 ax
11 x 11 x  11 ay
11 x 11 x  11 az

11 x 11 x  11 phitilde
11 x 11 Type:  11 function

11

IllinoisGRMHDxCon2Prim (conditional) x illinoisgrmhdxaxixouterxboundaries x apply linear extrapolation bcs on ax{mu}, so that bcs are flat on bx x

11 x 11 After:  11 illinoisgrmhdxsync
11 x 11 Language: 11 c
11 x 11 Reads:  11 phitilde

11 x 11 x  11 ax
11 x 11 x  11 ay
11 x 11 x  11 az

11 x 11 Type:  11 function
11 x 11 Writes: 11 phitilde(boundary)
11 x 11 x  11 ax(boundary)
11 x 11 x  11 ay(boundary)
11 x 11 x  11 az(boundary)

11

IllinoisGRMHDxCon2Prim (conditional) x illinoisgrmhdxcomputexbxandxbstaggerxfromxa x compute b and bxstagger from a sync: grmhdxprimitivesxbi,grmhdxprimitivesxbixstagger x

11 x 11 After:  11 illinoisgrmhdxaxixouterxboundaries
11 x 11 Language: 11 c
11 x 11 Reads:  11 grid::coordinates
11 x 11 x  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 phitilde

11 x 11 x  11 ax
11 x 11 x  11 ay
11 x 11 x  11 az

11 x 11 Sync:  11 grmhdxbxcenter
11 x 11 x  11 grmhdxbxstagger
11 x 11 Type:  11 function
11 x 11 Writes: 11 phitilde(everywhere)
11 x 11 x  11 ax(everywhere)
11 x 11 x  11 ay(everywhere)
11 x 11 x  11 az(everywhere)
11 x 11 x  11 grmhdxbxcenter(everywhere)
11 x 11 x  11 grmhdxbxstagger(everywhere)

11

IllinoisGRMHDxCon2Prim (conditional) x illinoisgrmhdxperturbxconservatives x perturb conservatives before con2prim x

11 x 11 After: 11 illinoisgrmhdxcomputexbxandxbstaggerxfromxa
11 x 11 Before: 11 illinoisgrmhdxconservsxtoxprims
11 x 11 Type: 11 group

11

IllinoisGRMHDxCon2Prim (conditional) x illinoisgrmhdxconservsxtoxprims x compute primitive variables from conservatives x

11 x 11 After: 11 illinoisgrmhdxcomputexbxandxbstaggerxfromxa
11 x 11 Type: 11 group

11

IllinoisGRMHDxCon2Prim (conditional) x illinoisgrmhdxhydroxouterxboundaries x apply selected boundary conditions to the primitive variables x

11 x 11 After: 11 illinoisgrmhdxconservsxtoxprims
11 x 11 Type: 11 group

11

AddToTmunu (conditional) x illinoisgrmhdxcomputextmunu x compute stress-energy tensor

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 grmhdxvelocities

11 x 11 x  11 u0

11 x 11 x  11 grmhdxbxcenter
11 x 11 Type:  11 function
11 x 11 Writes: 11 tmunubase::stressxenergyxscalar(everywhere)
11 x 11 x  11 tmunubase::stressxenergyxvector(everywhere)
11 x 11 x  11 tmunubase::stressxenergyxtensor(everywhere)

11

MoLxCalcRHS (conditional) x illinoisgrmhdxrhs x evaluate rhss grhd equations x

11 x 11 Type: 11 group

11

CCTKxBASEGRID (conditional) x illinoisgrmhdxinitsymbound x schedule symmetries x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

IllinoisGRMHDxRHS (conditional) x illinoisgrmhdxevaluatexsourcesxrhs x evaluate source terms in grmhd rhss x

11 x 11 Type: 11 group

11

IllinoisGRMHDxRHS (conditional) x illinoisgrmhdxevaluatexfluxesxrhs x evaluate flux terms in grmhd rhss x

11 x 11 After: 11 illinoisgrmhdxevaluatexsourcesxrhs
11 x 11 Type: 11 group

11

IllinoisGRMHDxRHS (conditional) x convertxillinoisgrmhdxtoxhydrobase x convert needed hydrobase variables for nrpyleakage

11 x 11 After:  11 illinoisgrmhdxevaluatexfluxesxrhs
11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 grmhdxvelocities
11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobase::vel(everywhere)

11

IllinoisGRMHDxRHS x illinoisgrmhdxevaluatexphitildexandxaxgaugexrhs x evaluate phitilde rhs and gauge contributions to axi rhs x

11 x 11 After:  11 evaluatexfluxesxrhs
11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 admbase::metric
11 x 11 x  11 phitilde

11 x 11 x  11 ax
11 x 11 x  11 ay
11 x 11 x  11 az

11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxprimitivesxreconstructedxtemps
11 x 11 x  11 emxrhs

11

CCTKxANALYSIS (conditional) x convertxillinoisgrmhdxtoxhydrobase x convert illinoisgrmhd-native variables to hydrobase x

11 x 11 After:  11 mlxbssnxevolcalcgroup
11 x 11 Before: 11 computexbixb2xpoynxfluxet
11 x 11 x  11 convertxtoxmhdx3velocity
11 x 11 x  11 particlextraceret
11 x 11 x  11 volumeintegralgroup
11 x 11 Language: 11 c
11 x 11 Options: 11 global-early
11 x 11 x  11 loop-local
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobase::vel(everywhere)
11 x 11 x  11 hydrobase::wxlorentz(everywhere)
11 x 11 x  11 hydrobase::bvec(everywhere)

11

IllinoisGRMHDxprimsxtoxconservs (conditional) x illinoisgrmhdxhybridxentropyxprimsxtoxconservs x entropy+hybrid version of illinoisgrmhdxprimsxtoxconservs

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 Sync:  11 grmhdxconservatives
11 x 11 x  11 entxstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::entropy(everywhere)

11

IllinoisGRMHDxconservsxtoxprims (conditional) x illinoisgrmhdxhybridxentropyxconservsxtoxprims x entropy+hybrid version of illinoisgrmhdxconservsxtoxprims x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grid::coordinates
11 x 11 x  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 grmhdxconservatives
11 x 11 x  11 entxstar
11 x 11 x  11 grmhdxbxcenter
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 failurexchecker(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::entropy(everywhere)

11

IllinoisGRMHDxhydroxouterxboundaries (conditional) x illinoisgrmhdxhybridxentropyxhydroxouterxboundaries x entropy+hybrid version of illinoisgrmhdxhydroxouterxboundaries

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 Sync:  11 grmhdxvelocities
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 hydrobase::entropy
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::entropy(everywhere)

11

IllinoisGRMHDxevaluatexsourcesxrhs (conditional) x illinoisgrmhdxhybridxentropyxevaluatexsourcesxrhs x entropy+hybrid version of illinoisgrmhdxevaluatexsourcesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 admbase::curv
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservativesxrhs
11 x 11 x  11 entxstarxrhs
11 x 11 x  11 emxrhs

11

IllinoisGRMHDxevaluatexfluxesxrhs (conditional) x illinoisgrmhdxhybridxentropyxevaluatexfluxesxrhs x entropy+hybrid version of illinoisgrmhdxevaluatexfluxesxrhs

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 x  11 grmhdxbxstagger
11 x 11 x  11 phitilde

11 x 11 x  11 ax
11 x 11 x  11 ay
11 x 11 x  11 az

11 x 11 x  11 grmhdxconservativesxrhs(interior)
11 x 11 x  11 entxstarxrhs(interior)
11 x 11 x  11 emxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxfluxxtemps
11 x 11 x  11 grmhdxconservativesxrhs
11 x 11 x  11 entxstarxflux
11 x 11 x  11 entxstarxrhs
11 x 11 x  11 grmhdxprimitivesxreconstructedxtemps
11 x 11 x  11 grmhdxcminxcmaxxtemps
11 x 11 x  11 emxrhs

11

HydroBasexPrim2ConInitial (conditional) x illinoisgrmhdxprim2con2prim x convert hydrobase initial data to illinoisgrmhd variables x

11 x 11 Type: 11 group

11

IllinoisGRMHDxperturbxprimitives (conditional) x illinoisgrmhdxhybridxentropyxperturbxprimitives x entropy+hybrid version of illinoisgrmhdxperturbxprimitives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 phitilde

11 x 11 x  11 ax
11 x 11 x  11 ay
11 x 11 x  11 az

11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 phitilde(everywhere)
11 x 11 x  11 ax(everywhere)
11 x 11 x  11 ay(everywhere)
11 x 11 x  11 az(everywhere)

11

IllinoisGRMHDxperturbxconservatives (conditional) x illinoisgrmhdxhybridxentropyxperturbxconservatives x entropy+hybrid version of illinoisgrmhdxperturbxconservatives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grmhdxconservatives
11 x 11 x  11 entxstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)

11

IllinoisGRMHDxprimsxtoxconservs (conditional) x illinoisgrmhdxhybridxprimsxtoxconservs x hybrid version of illinoisgrmhdxprimsxtoxconservs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 Sync:  11 grmhdxconservatives
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)

11

IllinoisGRMHDxconservsxtoxprims (conditional) x illinoisgrmhdxhybridxconservsxtoxprims x hybrid version of illinoisgrmhdxconservsxtoxprims x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grid::coordinates
11 x 11 x  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 grmhdxconservatives
11 x 11 x  11 grmhdxbxcenter
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 failurexchecker(everywhere)
11 x 11 x  11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)

11

IllinoisGRMHDxhydroxouterxboundaries (conditional) x illinoisgrmhdxhybridxhydroxouterxboundaries x hybrid version of illinoisgrmhdxouterxboundaries

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 Sync:  11 grmhdxvelocities
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)

11

IllinoisGRMHDxevaluatexsourcesxrhs (conditional) x illinoisgrmhdxhybridxevaluatexsourcesxrhs x hybrid version of illinoisgrmhdxevaluatexsourcesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 admbase::curv
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservativesxrhs
11 x 11 x  11 emxrhs

11

IllinoisGRMHDxevaluatexfluxesxrhs (conditional) x illinoisgrmhdxhybridxevaluatexfluxesxrhs x hybrid version of illinoisgrmhdxevaluatexfluxesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 x  11 grmhdxbxstagger
11 x 11 x  11 phitilde

11 x 11 x  11 ax
11 x 11 x  11 ay
11 x 11 x  11 az

11 x 11 x  11 grmhdxconservativesxrhs(interior)
11 x 11 x  11 emxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxfluxxtemps
11 x 11 x  11 grmhdxconservativesxrhs
11 x 11 x  11 grmhdxcminxcmaxxtemps
11 x 11 x  11 grmhdxprimitivesxreconstructedxtemps
11 x 11 x  11 emxrhs

IllinoisGRMHDxperturbxprimitives (conditional) x illinoisgrmhdxhybridxperturbxprimitives x hybrid version of illinoisgrmhdxperturbxprimitives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 phitilde

11 x 11 x  11 ax
11 x 11 x  11 ay
11 x 11 x  11 az

11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 phitilde(everywhere)
11 x 11 x  11 ax(everywhere)
11 x 11 x  11 ay(everywhere)
11 x 11 x  11 az(everywhere)

11

IllinoisGRMHDxperturbxconservatives (conditional) x illinoisgrmhdxhybridxperturbxconservatives x hybrid version of illinoisgrmhdxperturbxconservatives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grmhdxconservatives
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)

11

IllinoisGRMHDxprimsxtoxconservs (conditional) x illinoisgrmhdxtabulatedxentropyxprimsxtoxconservs x entropy+tabulated version of illinoisgrmhdxprimsxtoxconservs

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 Sync:  11 grmhdxconservatives
11 x 11 x  11 yexstar
11 x 11 x  11 entxstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 yexstar(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::entropy(everywhere)
11 x 11 x  11 hydrobase::yxe(everywhere)
11 x 11 x  11 hydrobase::temperature(everywhere)

11

IllinoisGRMHDxPrim2Con2Prim (conditional) x convertxhydrobasextoxillinoisgrmhd x convert hydrobase initial data (id) to illinoisgrmhd variables x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::vel
11 x 11 x  11 hydrobase::avec
11 x 11 x  11 hydrobase::aphi
11 x 11 Sync:  11 grmhdxvelocities

11 x 11 x  11 ax
11 x 11 x  11 ay
11 x 11 x  11 az

11 x 11 x  11 phitilde
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxvelocities(everywhere)
11 x 11 x  11 phitilde(everywhere)
11 x 11 x  11 ax(everywhere)
11 x 11 x  11 ay(everywhere)
11 x 11 x  11 az(everywhere)

11

IllinoisGRMHDxconservsxtoxprims (conditional) x illinoisgrmhdxtabulatedxentropyxconservsxtoxprims x entropy+tabulated version of illinoisgrmhdxconservsxtoxprims

11 x 11 Language: 11 c
11 x 11 Reads:  11 grid::coordinates
11 x 11 x  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 grmhdxconservatives
11 x 11 x  11 entxstar
11 x 11 x  11 yexstar
11 x 11 x  11 grmhdxbxcenter
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 yexstar(everywhere)
11 x 11 x  11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 failurexchecker(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::entropy(everywhere)
11 x 11 x  11 hydrobase::yxe(everywhere)
11 x 11 x  11 hydrobase::temperature(everywhere)

11

IllinoisGRMHDxhydroxouterxboundaries (conditional) x illinoisgrmhdxtabulatedxentropyxhydroxouterxboundaries x entropy+tabulated version of illinoisgrmhdxouterxboundaries x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 Sync:  11 grmhdxvelocities
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 yexstar(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::entropy(everywhere)
11 x 11 x  11 hydrobase::yxe(everywhere)
11 x 11 x  11 hydrobase::temperature(everywhere)

11

IllinoisGRMHDxevaluatexsourcesxrhs (conditional) x illinoisgrmhdxtabulatedxentropyxevaluatexsourcesxrhs x entropy+tabulated version of illinoisgrmhdxevaluatexsourcesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 admbase::curv
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservativesxrhs
11 x 11 x  11 entxstarxrhs
11 x 11 x  11 yexstarxrhs
11 x 11 x  11 emxrhs

11

IllinoisGRMHDxevaluatexfluxesxrhs (conditional) x illinoisgrmhdxtabulatedxentropyxevaluatexfluxesxrhs x entropy+tabulated version of illinoisgrmhdxevaluatexfluxesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 x  11 grmhdxbxstagger
11 x 11 x  11 phitilde

11 x 11 x  11 ax
11 x 11 x  11 ay
11 x 11 x  11 az

11 x 11 x  11 grmhdxconservativesxrhs(interior)
11 x 11 x  11 emxrhs(interior)
11 x 11 x  11 entxstarxrhs(interior)
11 x 11 x  11 yexstarxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxfluxxtemps
11 x 11 x  11 entxstarxflux
11 x 11 x  11 yexstarxflux
11 x 11 x  11 grmhdxconservativesxrhs
11 x 11 x  11 entxstarxrhs
11 x 11 x  11 yexstarxrhs
11 x 11 x  11 grmhdxcminxcmaxxtemps
11 x 11 x  11 grmhdxprimitivesxreconstructedxtemps
11 x 11 x  11 emxrhs

11

IllinoisGRMHDxperturbxprimitives (conditional) x illinoisgrmhdxtabulatedxentropyxperturbxprimitives x entropy+tabulated version of illinoisgrmhdxperturbxprimitives

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 phitilde

11 x 11 x  11 ax
11 x 11 x  11 ay
11 x 11 x  11 az

11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::entropy
11 x 11 x  11 hydrobase::yxe(everywhere)
11 x 11 x  11 hydrobase::temperature(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 phitilde(everywhere)
11 x 11 x  11 ax(everywhere)
11 x 11 x  11 ay(everywhere)
11 x 11 x  11 az(everywhere)

11

IllinoisGRMHDxperturbxconservatives (conditional) x illinoisgrmhdxtabulatedxentropyxperturbxconservatives x entropy+tabulated version of illinoisgrmhdxperturbxconservatives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grmhdxconservatives
11 x 11 x  11 entxstar
11 x 11 x  11 yexstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 entxstar(everywhere)
11 x 11 x  11 yexstar(everywhere)

11

IllinoisGRMHDxprimsxtoxconservs (conditional) x illinoisgrmhdxtabulatedxprimsxtoxconservs x tabulated version of illinoisgrmhdxprimsxtoxconservs

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 Sync:  11 grmhdxconservatives
11 x 11 x  11 yexstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 yexstar(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::yxe(everywhere)
11 x 11 x  11 hydrobase::temperature(everywhere)

11

IllinoisGRMHDxconservsxtoxprims (conditional) x illinoisgrmhdxtabulatedxconservsxtoxprims x tabulated version of illinoisgrmhdxconservsxtoxprims x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grid::coordinates
11 x 11 x  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 grmhdxconservatives
11 x 11 x  11 yexstar
11 x 11 x  11 grmhdxbxcenter
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 yexstar(everywhere)
11 x 11 x  11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 failurexchecker(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::yxe(everywhere)
11 x 11 x  11 hydrobase::temperature(everywhere)

11

IllinoisGRMHDxhydroxouterxboundaries (conditional) x illinoisgrmhdxtabulatedxhydroxouterxboundaries x tabulated version of illinoisgrmhdxouterxboundaries

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 Sync:  11 grmhdxvelocities
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 yexstar(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::yxe(everywhere)
11 x 11 x  11 hydrobase::temperature(everywhere)

11

IllinoisGRMHDxevaluatexsourcesxrhs (conditional) x illinoisgrmhdxtabulatedxevaluatexsourcesxrhs x tabulated version of illinoisgrmhdxevaluatexsourcesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 admbase::curv
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservativesxrhs
11 x 11 x  11 yexstarxrhs
11 x 11 x  11 emxrhs

11

IllinoisGRMHDxPrim2Con2Prim (conditional) x illinoisgrmhdxperturbxprimitives x perturb initial primitive data x

11 x 11 After: 11 convertxhydrobasextoxillinoisgrmhd
11 x 11 Before: 11 illinoisgrmhdxprimsxtoxconservs
11 x 11 Type: 11 group

11

IllinoisGRMHDxevaluatexfluxesxrhs (conditional) x illinoisgrmhdxtabulatedxevaluatexfluxesxrhs x tabulated version of illinoisgrmhdxevaluatexfluxesxrhs

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 x  11 grmhdxbxstagger
11 x 11 x  11 phitilde

11 x 11 x  11 ax
11 x 11 x  11 ay
11 x 11 x  11 az

11 x 11 x  11 grmhdxconservativesxrhs(interior)
11 x 11 x  11 yexstarxrhs(interior)
11 x 11 x  11 emxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxfluxxtemps
11 x 11 x  11 grmhdxconservativesxrhs
11 x 11 x  11 yexstarxflux
11 x 11 x  11 yexstarxrhs
11 x 11 x  11 grmhdxprimitivesxreconstructedxtemps
11 x 11 x  11 grmhdxcminxcmaxxtemps
11 x 11 x  11 emxrhs

11

IllinoisGRMHDxperturbxprimitives (conditional) x illinoisgrmhdxtabulatedxperturbxprimitives x tabulated version of illinoisgrmhdxperturbxprimitives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 hydrobase::yxe
11 x 11 x  11 hydrobase::temperature
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 phitilde

11 x 11 x  11 ax
11 x 11 x  11 ay
11 x 11 x  11 az

11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxvelocities(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 hydrobase::yxe(everywhere)
11 x 11 x  11 hydrobase::temperature(everywhere)
11 x 11 x  11 phitilde(everywhere)
11 x 11 x  11 ax(everywhere)
11 x 11 x  11 ay(everywhere)
11 x 11 x  11 az(everywhere)

11

IllinoisGRMHDxperturbxconservatives (conditional) x illinoisgrmhdxtabulatedxperturbxconservatives x tabulated version of illinoisgrmhdxperturbxconservatives

11 x 11 Language: 11 c
11 x 11 Reads:  11 grmhdxconservatives
11 x 11 x  11 yexstar
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 yexstar(everywhere)

11

CCTKxWRAGH (conditional) x illinoisgrmhdxbackwardxcompatiblexinitialize x set up the grhayl structs and read the table for tabulated eos. x

11 x 11 Language: 11 c
11 x 11 Options: 11 global
11 x 11 Type:  11 function

11

CCTKxPOSTINITIAL (conditional) x illinoisgrmhdxbackwardxcompatiblexdata x copy data to deprecated variables x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 phitilde
11 x 11 x  11 grmhdxbxcenter
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxprimitivesxallbutbi
11 x 11 x  11 psi6phi
11 x 11 x  11 grmhdxprimitivesxbi

11

MoLxPostRHS (conditional) x illinoisgrmhdxbackwardxcompatiblexdata x copy data to deprecated variables x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 phitilde
11 x 11 x  11 grmhdxbxcenter
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxprimitivesxallbutbi
11 x 11 x  11 psi6phi
11 x 11 x  11 grmhdxprimitivesxbi

11

IllinoisGRMHDxPrim2Con2Prim (conditional) x convertxillinoisgrmhdxtoxhydrobase x convert illinoisgrmhd-native variables to hydrobase

11 x 11 After:  11 illinoisgrmhdxconservsxtoxprims
11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobase::vel(everywhere)
11 x 11 x  11 hydrobase::wxlorentz(everywhere)
11 x 11 x  11 hydrobase::bvec(everywhere)

11

CCTKxANALYSIS (conditional) x convertxillinoisgrmhdxtoxhydrobase x convert illinoisgrmhd-native variables to hydrobase x

11 x 11 After:  11 mlxbssnxevolcalcgroup
11 x 11 Before: 11 computexbixb2xpoynxfluxet
11 x 11 x  11 convertxtoxmhdx3velocity
11 x 11 x  11 particlextraceret
11 x 11 x  11 volumeintegralgroup
11 x 11 Language: 11 c
11 x 11 Options: 11 global-early
11 x 11 x  11 loop-local
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobase::vel(everywhere)
11 x 11 x  11 hydrobase::wxlorentz(everywhere)
11 x 11 x  11 hydrobase::bvec(everywhere)

11

IllinoisGRMHDxprimsxtoxconservs (conditional) x illinoisgrmhdxhybridxprimsxtoxconservs x hybrid version of illinoisgrmhdxprimsxtoxconservs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 grmhdxbxcenter
11 x 11 Sync:  11 grmhdxconservatives
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)

11

IllinoisGRMHDxconservsxtoxprims (conditional) x illinoisgrmhdxhybridxconservsxtoxprims x hybrid version of illinoisgrmhdxconservsxtoxprims x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grid::coordinates
11 x 11 x  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 grmhdxconservatives
11 x 11 x  11 grmhdxbxcenter
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)
11 x 11 x  11 failurexchecker(everywhere)
11 x 11 x  11 u0(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)

11

IllinoisGRMHDxPrim2Con2Prim (conditional) x illinoisgrmhdxcomputexbxandxbstaggerxfromxa x compute b and bxstagger from a x

11 x 11 After:  11 convertxhydrobasextoxillinoisgrmhd
11 x 11 Language: 11 c
11 x 11 Reads:  11 grid::coordinates
11 x 11 x  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 phitilde

11 x 11 x  11 ax
11 x 11 x  11 ay
11 x 11 x  11 az

11 x 11 Sync:  11 grmhdxbxcenter
11 x 11 x  11 grmhdxbxstagger
11 x 11 Type:  11 function
11 x 11 Writes: 11 phitilde(everywhere)
11 x 11 x  11 ax(everywhere)
11 x 11 x  11 ay(everywhere)
11 x 11 x  11 az(everywhere)
11 x 11 x  11 grmhdxbxcenter(everywhere)
11 x 11 x  11 grmhdxbxstagger(everywhere)

11

IllinoisGRMHDxhydroxouterxboundaries (conditional) x illinoisgrmhdxhybridxhydroxouterxboundaries x hybrid version of illinoisgrmhdxouterxboundaries

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 Sync:  11 grmhdxvelocities
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 Type:  11 function
11 x 11 Writes: 11 u0(everywhere)
11 x 11 x  11 grmhdxconservatives(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)

11

IllinoisGRMHDxevaluatexsourcesxrhs (conditional) x illinoisgrmhdxhybridxevaluatexsourcesxrhs x hybrid version of illinoisgrmhdxevaluatexsourcesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 admbase::curv
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservativesxrhs
11 x 11 x  11 emxrhs

11

IllinoisGRMHDxevaluatexfluxesxrhs (conditional) x illinoisgrmhdxhybridxevaluatexfluxesxrhs x hybrid version of illinoisgrmhdxevaluatexfluxesxrhs x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 x  11 grmhdxbxstagger
11 x 11 x  11 phitilde

11 x 11 x  11 ax
11 x 11 x  11 ay
11 x 11 x  11 az

11 x 11 x  11 grmhdxconservativesxrhs(interior)
11 x 11 x  11 emxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxfluxxtemps
11 x 11 x  11 grmhdxconservativesxrhs
11 x 11 x  11 grmhdxcminxcmaxxtemps
11 x 11 x  11 grmhdxprimitivesxreconstructedxtemps
11 x 11 x  11 emxrhs

IllinoisGRMHDxperturbxprimitives (conditional) x illinoisgrmhdxhybridxperturbxprimitives x hybrid version of illinoisgrmhdxperturbxprimitives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 hydrobase::rho
11 x 11 x  11 hydrobase::press
11 x 11 x  11 hydrobase::eps
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 phitilde

11 x 11 x  11 ax
11 x 11 x  11 ay
11 x 11 x  11 az

11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobase::rho(everywhere)
11 x 11 x  11 hydrobase::press(everywhere)
11 x 11 x  11 hydrobase::eps(everywhere)
11 x 11 x  11 grmhdxvelocities(everywhere)
11 x 11 x  11 phitilde(everywhere)
11 x 11 x  11 ax(everywhere)
11 x 11 x  11 ay(everywhere)
11 x 11 x  11 az(everywhere)

11

IllinoisGRMHDxperturbxconservatives (conditional) x illinoisgrmhdxhybridxperturbxconservatives x hybrid version of illinoisgrmhdxperturbxconservatives x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grmhdxconservatives
11 x 11 Type:  11 function
11 x 11 Writes: 11 grmhdxconservatives(everywhere)

11

IllinoisGRMHDxPrim2Con2Prim (conditional) x illinoisgrmhdxprimsxtoxconservs x compute conservative variables from primitives x

11 x 11 After: 11 illinoisgrmhdxcomputexbxandxbstaggerxfromxa
11 x 11 Type: 11 group

11

IllinoisGRMHDxPrim2Con2Prim (conditional) x illinoisgrmhdxconservsxtoxprims x compute primitive variables from conservatives x

11 x 11 After: 11 illinoisgrmhdxprimsxtoxconservs
11 x 11 Type: 11 group

11

IllinoisGRMHDxPrim2Con2Prim (conditional) x convertxillinoisgrmhdxtoxhydrobase x convert illinoisgrmhd-native variables to hydrobase x

11 x 11 After:  11 illinoisgrmhdxconservsxtoxprims
11 x 11 Language: 11 c
11 x 11 Reads:  11 admbase::metric
11 x 11 x  11 admbase::lapse
11 x 11 x  11 admbase::shift
11 x 11 x  11 grmhdxvelocities
11 x 11 x  11 grmhdxbxcenter
11 x 11 Type:  11 function
11 x 11 Writes: 11 hydrobase::vel(everywhere)
11 x 11 x  11 hydrobase::wxlorentz(everywhere)
11 x 11 x  11 hydrobase::bvec(everywhere)

11

Aliased Functions x

11 Alias Name: xxxxxxx 11 Function Name:
11 IllinoisGRMHDxsync 11 IllinoisGRMHDxsyncxconservatives